# supplement_jasa

### v1

@@ Abstract
Building artificially intelligent geospatial systems require rapid delivery of spatial data analysis at massive scales with minimal human intervention. Depending upon their intended use, data analysis may also entail model assessment and uncertainty quantification. This article devises transfer learning frameworks for deployment in artificially intelligent systems, where a massive data set is split into smaller data sets that stream into the analytical framework to propagate learning and assimilate inference for the entire data set. Specifically, we introduce Bayesian predictive stacking for multivariate spatial data and demonstrate its effectiveness in rapidly analyzing massive data sets. Furthermore, we make inference feasible in a reasonable amount of time, and without excessively demanding hardware settings. We illustrate the effectiveness of this approach in extensive simulation experiments and subsequently analyze massive data sets in climate science on sea surface temperatures and on vegetation index.
@@ 1 Introduction
To analyze massive amounts of data to be accessed in a GeoAI system, we devise Bayesian transfer learning over a streaming set of spatial datasets. We envisage a setting where the data stream into the GeoAI system in the form of $K$ subsets, where each subset is composed of $\approx n/K$ locations with $K$ being a function of computational resources available to GeoAI. We analyze and transfer learning from one dataset to the next and assimilate inference from these subsets to the full data set over $n$ locations. The concept of delegating Bayesian calculations to a group of independent datasets is intuitive and has been explored from diverse perspectives. Examples include Consensus Monte Carlo (cmc) (Scott et al.,, 2016) and its kernel-based adaptations (Rendell et al.,, 2021), Bayesian meta-analysis in clinical applications (Parmigiani,, 2002, see, e.g., Chapter 4 in) and in spatially-temporally structured data (Bell et al.,, 2005; Kang et al.,, 2011); and, more recently, approaches such as “meta-kriging” using geometric medians of posterior distributions (Srivastava et al.,, 2015; Minsker et al.,, 2017; Guhaniyogi and Banerjee,, 2018, 2019; Guhaniyogi et al.,, 2019). However, each of these approaches have, hitherto, relied upon MCMC or other iterative algorithms for estimating the weakly identifiable parameter. We, instead, propose to execute transfer learning using a double stacking approach. We first obtain stacked inference for each of the parameters within each subset and, then, assimilate inference across subsets by a second stacking algorithm. A flowchart is supplied in Figure 1 comprising three steps: (i) partition the data into smaller subsets; (ii) analyze each subset by stacking conjugate posteriors in closed form; and (iii) stack across data sets to deliver inference for the entire data.
@@
The contents of the paper evolve in accordance with Figure 1. Section 2 outlines Bayesian transfer learning using double bps. Section 3 provides details on computational aspects. Section 4 presents simulation experiments designed for empirically evaluating double bps. Section 5 analyzes a very large climate data set with observed locations in the millions at global scales. Finally, Section 6 concludes the paper with some discussion.
@@ 2.3 Accelerated learning for spatial random fields
For each dataset $\mathscr{D}_{k}$, bps computes: (i) an estimate of the posterior predictive $\hat{p}(\cdot\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$ for $k=1,\dots,K$ and $j=1,\dots,J$; and (ii) a set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$. Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\cdot\mid\mathscr{D}_{k})$ over the $k$ subsets. This double bps of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{k,i}\mid\mathscr{D}_{k,[-l]}\right)$ Once the stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are obtained, the double bps estimation of any posterior or posterior predictive distributions of interest is obtained from
@@ 6 Discussion
This manuscript lays out a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to handle massive spatial datasets with minimal human intervention. The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference on multivariate processes and circumventing iterative algorithms that require, often extensive, human tuning or control. Our proposed GeoSpatial AI system is capable of handling truly massive amounts of spatial data that preclude accessibility even on high performance CPU architectures and we develop Bayesian transfer learning using double Bayesian predictive stacking.

### v2

@@ Abstract
Building artificially intelligent geospatial systems require rapid delivery of spatial data analysis at massive scales with minimal human intervention. Depending upon their intended use, data analysis may also entail model assessment and uncertainty quantification. This article devises transfer learning frameworks for deployment in artificially intelligent systems, where a massive data set is split into smaller data sets that stream into the analytical framework to propagate learning and assimilate inference for the entire data set. Specifically, we introduce Bayesian predictive stacking for multivariate spatial data and demonstrate rapid and automated analysis of massive data sets. Furthermore, inference is delivered without human intervention without excessively demanding hardware settings. We illustrate the effectiveness of our approach through extensive simulation experiments and in producing inference from massive datasets on sea surface temperatures and vegetation index that are indistinguishable from traditional (and more expensive) statistical approaches.
@@ 1 Introduction
To analyze massive amounts of data to be accessed in a GeoAI system, we devise Bayesian transfer learning over a streaming set of spatial datasets. We envisage a setting where the data stream into the GeoAI system in the form of $K$ subsets, where each subset is composed of $\approx n/K$ locations with $K$ being a function of computational resources available to GeoAI. We analyze and transfer learning from one dataset to the next and assimilate inference from these subsets to the full data set over $n$ locations. The concept of delegating Bayesian calculations to a group of independent datasets is intuitive and has been explored from diverse perspectives. Examples include Consensus Monte Carlo (cmc) (Scott et al.,, 2016) and its kernel-based adaptations (Rendell et al.,, 2021), Bayesian meta-analysis in clinical applications (Parmigiani,, 2002, see, e.g., Chapter 4 in) and in spatially-temporally structured data (Bell et al.,, 2005; Kang et al.,, 2011); and, more recently, approaches such as “meta-kriging” using geometric medians of posterior distributions (Srivastava et al.,, 2015; Minsker et al.,, 2017; Guhaniyogi and Banerjee,, 2018, 2019; Guhaniyogi et al.,, 2019). However, each of these approaches have, hitherto, relied upon MCMC or other iterative algorithms for estimating the weakly identifiable parameter. We, instead, propose to execute transfer learning using a double stacking approach. We first obtain stacked inference for each of the parameters within each subset and, then, assimilate inference across subsets by a second stacking algorithm. A flowchart is supplied in Figure 1 comprising three steps: (i) partition data into smaller subsets; (ii) analyze each subset by stacking conjugate posteriors in closed form; and (iii) stack across data sets to deliver inference for the entire data.
@@
The contents of the paper evolve in accordance with Figure 1. Section 2 outlines Bayesian transfer learning using double bps. Section 3 provides details on computational aspects. Section 4 presents simulation experiments designed for empirically evaluating double bps. Section 5 analyzes a very large climate data set with observed locations in the millions at global scales. Finally, Section 6 concludes with some discussion. References to an accompanying Supplement are provided throughout the article and after Section 6.
@@ 2.3 Accelerated learning for spatial random fields
For each dataset $\mathscr{D}_{k}$, bps computes: (i) an estimate of the posterior predictive $\hat{p}(\cdot\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$ for $k=1,\dots,K$ and $j=1,\dots,J$; and (ii) a set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$. Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\cdot\mid\mathscr{D}_{k})$ over the $k$ subsets. This double bps of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{i}\mid\mathscr{D}_{k}\right)$ (see Section LABEL:sec:BPSobjfun). Once the stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are obtained, the double bps estimation of any posterior or posterior predictive distributions of interest is obtained from
@@ 6 Discussion
This manuscript devises a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to analyze massive datasets with minimal human intervention. The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference on multivariate processes and circumventing iterative algorithms that require, often extensive, human tuning or control. Our proposed GeoSpatial AI system is capable of handling truly massive amounts of spatial data that preclude accessibility even on high performance CPU architectures and we develop Bayesian transfer learning using double Bayesian predictive stacking.